### v1

# Can Large Language Models Act as Ensembler for Multi-GNNs?

Abstract.
Graph Neural Networks (GNNs) have emerged as powerful models for learning from graph-structured data. However, GNNs lack the inherent semantic understanding capability of rich textual nodes attributes, limiting their effectiveness in applications. On the other hand, we empirically observe that for existing GNN models, no one can consistently outperforms others across diverse datasets. In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. The model first aligns multiple GNNs, mapping the representations of different GNNs into the same space. Then, through LoRA fine-tuning, it aligns the space between the GNN and the LLM, injecting graph tokens and textual information into LLMs. This allows LensGNN to ensemble multiple GNNs and leverage LLM’s strengths, leading to a deeper understanding of both textual semantic information and graph structural information. Experimental results show that LensGNN  outperforms existing models. This research advances text-attributed graph ensemble learning by providing a robust, superior solution for integrating semantic and structural information. We provide our code and data here: https://anonymous.4open.science/r/EnsemGNN-E267/.
Large Language Models, Text-attributed Graph, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. INTRODUCTION
Graphs are structured data that captures the inter-relations between entities in the real world. To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced (Kipf and Welling, 2016) and message passing is the main mechanism for neighborhood aggregation (Hamilton et al., 2017a; Veličković et al., 2017). GNNs have demonstrated significant success across various applications such as social network analysis (Hamilton et al., 2017a), recommendation systems (Wang et al., 2019), and molecular property prediction (Gilmer et al., 2017).
Despite the success, two major challenges remain unresolved. First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes (Zhao et al., 2019). This can lead to the loss of valuable semantic information during the learning process, limiting the effectiveness of GNNs in applications where node features contain meaningful texts, such as citation networks and social media platforms. Second, while there have been a series of GNNs proposed, no one has been shown to consistently lead others across datasets of various domains. For example, we compare the node classification performance of four representative GNNs, namely, APPNP (Gasteiger et al., 2018), GAT (Veličković et al., 2017), GCN (Kipf and Welling, 2016) and GIN (Xu et al., 2019) on three benchmark datasets (Yan et al., 2023): Cora, Citeseer and PubMed. For fairness, we set the equal number of hidden representation layers and also the same dimensionality size. The results given in Figure 1 demonstrate that the winner varies across the datasets. The challenge of selecting the optimal GNN for a given dataset remains unresolved, as different GNN architectures exhibit varying strengths. This also restricts the wide applicability of GNNs.
[…]
Comparison of node classification accuracy for various GNN models across different datasets. To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
Figure 1. Comparison of node classification accuracy for various GNN models across different datasets. To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
To remedy the incapability of GNNs in semantic understanding, existing works (Qin et al., 2024; Li et al., 2024; Wang et al., 2024; Ye et al., 2024; Fang et al., 2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation. The inherent advantage of GNNs in utilizing graph structure, fused with the strength of LLMs in semantic understanding, mutually enhances both models and leads to superior performance in downstream tasks. However, how to select the optimal GNN in different scenarios remains a gap. Since the effectiveness of GNNs varies across datasets, there naturally arises a question: Can we develop an ensembled model for multi-GNNs? The ensembled model is expected to integrate the strengths of multiple GNNs and can consistently perform well across datasets. To further leverage the power of LLMs in text understanding, we thus upgrade the question: Can LLMs act as ensembler for multi-GNNs?
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. The model captures semantic information from node textual attributes using LLMs, while simultaneously enabling LLMs to implicitly evaluate and combine the strengths of multiple GNNs. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
  • We propose a novel method LensGNN to ensemble multi-GNNs with LLMs. To our best knowledge, we are the first work to attempt the trial. LensGNN combines the strengths of both GNNs and LLMs, which can be used to avoid the intricate process of manual GNN selection on various datasets.
  • We introduce an effective graph prompt tuning paradigm, which enhances LLMs’ capability in comprehending graph topology with two phases of alignment. The multi-GNN alignment maps multi-GNNs into the same low-dimensional space while the GNN-LLM alignment injects graph token representations into LLMs.
  • We conduct extensive experiments to show the effectiveness of LensGNN. In particular, we compare LensGNN with 17 state-of-the-art competitors. Our experimental results show that LensGNN outperforms baselines in a majority of cases, demonstrating its superior performance.
2. RELATED WORK
GNNs (Kipf and Welling, 2016; Xu et al., 2019; Veličković et al., 2017; Hamilton et al., 2017b) have been widely used in learning from graph-structured data. Existing GNN studies have been conducted on various types of graphs such as heterophilic graphs (Zhu et al., 2020; Pei et al., 2020; Bi et al., 2024), signed graphs (Derr et al., 2018; Huang et al., 2019) and temporal graphs (Rossi et al., 2020; Longa et al., 2023; Xu et al., 2020). Further, to improve the effectiveness of GNNs, there are also works (Nguyen et al., 2023; Rusch et al., 2023; Black et al., 2023) devoted to solving the notorious over-smoothing (Chen et al., 2020) and over-quashing problems (Akansha, 2023). In parallel, to scale GNNs to large-scale graphs, some works (Liao et al., 2022) focus on enhancing the model efficiency. A comprehensive GNN-related survey can be found at (Wu et al., 2020). Despite their advantages, GNNs struggle to process textual attributes effectively, as traditional GNNs lack the semantic understanding capability that is necessary to handle text-attributed graphs (TAGs). In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) (Zoph and Le, 2017). However, it is very computationally expensive, which necessitates new exploration.
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies (Jin et al., 2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, where models like GraphLLM (Chai et al., 2023) utilizes few-shot learning to address graph reasoning challenges. ENGINE (Zhu et al., 2024) is designed to efficiently integrate textual and topological information by adding a tunable lightweight GNN side structure alongside each layer of LLMs.Similarly, heuristic reasoning approaches, including Chain-of-Thought (CoT) (Wei et al., 2022) and models like StructGPT (Jiang et al., 2023), have demonstrated improved graph-based inference capabilities. Another notable category is to employ LLMs as encoders, as seen in the works of TextGNN (Zhu et al., 2021) and AdsGNN (Li et al., 2021), which focus on optimization strategies for encoding graph data. Employing LLMs as encoders involves using large language models to extract textual features, which are then used as initial feature vectors for nodes or edges in a graph. These feature vectors are passed into graph neural networks (GNNs) for further processing, allowing the model to incorporate both the textual information from the LLM and the structural information from the graph. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods such as LTRN (Zhang et al., 2021) and GLEM (Zhao et al., 2022), where iterative training helps improve cross-modality learning. LLMs and GNNs are aligned through iterative pseudo-label training (prediction alignment) or by using contrastive learning to align their latent representations (latent space alignment). Finally, graph-empowered LLMs like GreaseLM (Zhang et al., 2022) and DRAGON (Yasunaga et al., 2022) illustrate the potential of enhancing language models with graph-based knowledge to enrich semantic understanding. They integrate the strengths of large language models and graph structures by modifying Transformer architectures to encode text and graph data simultaneously, allowing for a more comprehensive understanding of the interconnected information within graphs. These diverse approaches highlight the evolving landscape of LLM and GNN integration. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
3. PRELIMINARIES
Text-attributed graph. A text-attributed graph is a graph where nodes and edges have textual attributes. In this work, we only consider the case where only nodes have textual attributes. Formally, a text-attributed graph can be defined as $G=(V,E,X^{(v)})$, where $V=\{v_{1},v_{2},\ldots,v_{|V|}\}$ represents the set of nodes, and $E=\{e_{1},e_{2},\ldots,e_{|E|}\}$ represents the set of edges, with $|V|=N$ indicating the total number of nodes in the graph. $X^{(v)}=\{x^{(v)}_{1},x^{(v)}_{2},\ldots,x^{(v)}_{|V|}\}$ denote the attributes on nodes, respectively, which are strings. It can be represented as $G=(V,E,\{x_{n}\}_{n\in V})$. Additionally, we define $A$ as the adjacency matrix of graph $G$, where the matrix size is $N\times N$. In an unweighted graph, $A(i,j)=1$ indicates that there is an edge between nodes $v_{i}$ and $v_{j}$, while $A(i,j)=0$ indicates that there is no edge between them.
Graph Neural Networks. Traditional GNNs are a class of deep learning models designed for handling graph-structured data. The basic architecture of a GNN includes a node representation layer and a message-passing layer. In the node representation layer, each node $v$ is assigned a feature vector $x_{v}$. In the message-passing layer, the representation vector $h_{v}^{(t)}$ of node $v$ is updated after the $t$-th iteration using the following formula:
(1) $h_{v}^{(t)}=\text{UPDATE}\left(\text{AGGREGATE}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$
where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$. The AGGREGATE function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state. Through this iterative process, GNNs are able to learn increasingly rich node representations, capturing both local and global structural information in the graph. The specific implementations of the AGGREGATE and UPDATE functions can vary depending on the particular GNN model.
4. METHODOLOGY
To enable the LLM to integrate multiple GNNs, alignment is primarily required in two aspects. First, the output representations of all the GNNs need to be aligned into a common vector space. Second, there must be alignment between the output representations of the GNNs and the semantic space required by the LLM. Based on this idea, we have designed the model in the following two steps. The structure of the whole model is shown in Fig. 2.
4.1. Aligning multi-GNNs
Given multiple GNNs, they could generate node representations in different low-dimensional spaces. Therefore, before ensembling GNNs, we need to first align them.
Figure 2. The overall process of LensGNN. Step 1 is the workflow of the pre-training phase of LensGNN’s GNNs. We employ a language pre-training encoder to extract semantic information from the textual attributes of nodes. Concurrently, we train the GNN encoder to capture structural information. During this phase, a MLP is utilized as a classifier, and the model is trained through gradient feedback. Step 2 is the workflow of the LoRA (Low-Rank Adaptation) supervised fine-tuning stage of LensGNN’s LLM. At this stage, we freeze the parameters of the GNNs and infer the hidden representations of the graph nodes. We segment GNN representations and convert them into graph tokens that LLM can read. The graph tokens, node text tokens, and other prompts are concatenated to form an instruction, which is then input into the LLMs for LoRA supervised fine-tuning.
[…]
In the pre-training phase, LensGNN consists of the following architecture: a pre-trained language encoder with frozen parameters, a GNN layer, a linear mapping layer, and an MLP classifier, all of which have parameters that are updated during training.In the fine-tuning phase of LensGNN, the parameters of the pre-trained language encoder and the GNN layers are frozen, and only the Lora parameters of the LLM are updated. In addition, the size of Hidden Representation is (1, t * e), where e is the length of the enbedding of each token in LLM, and t is the number of tokens each node representation is converted into.
Node feature initialization. To utilize the rich semantic information of textual features, we use pre-trained language models to initialize them. Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT (Reimers, 2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
(2) $\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$
This process allows texts of any dataset and length to be converted into the same feature space.
GNN representations. Next, the model takes the text representations obtained from Eq. 2 for each node and feeds them into multiple GNNs. For the $k$-th GNN, we denote the node embedding matrix $H^{[k]}$ generated by the final layer as:
(3) $H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$
where $\tilde{X}$ is the initial node embedding matrix from Eq. 2 and $\Theta_{G_{k}}$ is the learnable parameters of the $k$-th GNN. For each node $v_{i}$, let $h_{i}^{[k]}$ be its embedding vector, which is the $i$-th row in $H^{[k]}$.
Reshape node representations. The output of each GNN is then connected to a linear layer, which transforms node representations to the dimensionality of hidden embeddings in LLMs. This step paves the way for the subsequent GNN and LLM alignment. Details will be given in the next section. For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
(4) $\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$
Note that the shape of $\tilde{h}^{[k]}_{i}$ is a one-dimensional vector of length $t\times e$, where $t$ is a hyperparameter indicating the number of graph tokens each node representation is mapped to. The dimensionality $e$ comes from the hidden embedding layer of the LLM used, where each token is mapped to an embedding of shape $(1,e)$.
Multi-GNN alignment. To align node representations from multiple GNNs, we next feed them into a shared linear layer, which serves as the classifier. During the training time, we alternatively input representations from different GNNs into the classifier and train them sequentially. This training approach allows node representations from multi-GNNs to be aligned, which facilitates the integration of multi-GNNs. After training, the parameters of GNNs and the linear layer in Eq. 4 will be frozen, which are then used in aligning GNNs and LLMs. The overall procedure for GNN alignment is summarized in the top of Fig. 2.
In summary, during the multi-GNN alignment step, we train each GNN to ensure that their outputs reside in the same vector space. By incorporating the node feature initialization step that utilizes language pre-trained models, our model gains the ability to extract semantic and structural information from the nodes in the text graph. Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the integration of extensive graph structural information into the LLM in the subsequent steps.
4.2. Ensembling multi-GNNs with LLM
After GNNs are trained, we can get node representations capturing rich graph structural information. Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure (Guo et al., 2023). Therefore, in the second stage, we empower LLMs to comprehend graph representations. Further, with specially designed prompts, we enable these models to implicitly perform ensembling.
Align GNNs and LLMs. The key step in enabling LLMs to understand GNN tokens lies in aligning GNN representations, which encapsulate rich structural information, with the semantic space required by LLMs. This process necessitates fine-tuning LLMs. However, due to the extensive number of parameters in these models, fine-tuning requires significant computational resources, which is challenging for practical applications. As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector (Tang et al., 2024), mapping GNN representations into the semantic space. While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
[…]
Figure 3. A prompt template. The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific node text inputs and their multiple GNN representations.
Prompt design and Multi-GNN ensembling. After aligning graph structure and semantics, LLM can understand graph representations. We next design prompts and take LLM as the ensembler to integrate multi-GNNs. Our goal is to leverage the strengths of both multi-GNNs and LLM to output more accurate predictions. The prompt design follows three key principles:
  1. (1)
    Simultaneous input of textual strings from a node and its adjacent neighbors, graph tokens of the node, and task specification. Existing studies (Kipf and Welling, 2016; Veličković et al., 2017; Verma et al., 2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction. Further, clear task instructions are necessary for accurate prediction.
  2. (2)
    Differentiation between text tokens and graph tokens. It directs LLM to correctly distinguish between textual tokens and graph tokens within different segments of the prompt, thus preventing confusion.
  3. (3)
    Guidance for learning from multi-GNNs: It leads LLM to implicitly learn how to effectively combine strengths of different GNNs.
Based on the above principles, we provide a prompt template used in our experiments, as shown in Figure 3. Details on the prompt are given in the caption of the figure. In practice, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
5. EXPERIMENT
In this section, we conduct extensive experiments to evaluate the performance of our model across various datasets, aiming to answer the following questions:
  • RQ1: How does LensGNN perform compared to SOTA baseline models?
  • RQ2: What is the importance of different components in LensGNN?
  • RQ3: Can small language models serve as multi-GNN ensembler?
  • RQ4: How efficient is our method LensGNN ?
  • RQ5: Is our method LensGNN insensitive to model hyper-parameters?
Table 1. Datasets statistics.
Dataset # Nodes # Edges # Classes
Cora 2,708 5,429 7
PubMed 19,717 44,338 3
ogbn-arXiv 169,343 1,166,243 40
Table 2. Hyperparameter Search Spaces.
Hyperparameter Search space
learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
dropout $[0,0.25]$
loraplus lr ratio $\{16.0,24.0,32.0\}$
Graph token $t$$\{1,2,4,8,16\}$
Table 3. The classification performance over the methods on 3 datasets. We highlight the best score on each dataset in bold and underline the runner-up’s. The - means that we did not obtain results from their original paper.
Model type Model Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
MLP MLP 66.42 82.40 61.47
GNN GCN 85.97 83.78 70.81
GAT 86.71 83.59 70.05
GIN 85.60 82.03 67.03
GraphSAGE 84.87 87.01 70.35
Graphormer 80.41 88.75 72.81
GT 86.42 88.24 -
CoarFormer 88.69 89.75 71.66
Language Model BERT 80.15 93.91 72.78
SentenceBERT 78.82 92.49 71.42
DeBERTa 77.79 93.45 72.90
LLM Based DGTL 81.10 87.10 -
SNS-GPT4 82.50 93.80 74.40
InstructGraph-INS 89.33 81.09 -
InstructGLM 90.77 94.62 75.70
GAugLLM - 83.68 74.15
Baichuan2-13B 13.65 36.04 4.79
LensGNN-[GCN+GAT] 88.56 94.11 75.78
LensGNN-[GCN+GIN] 91.88 94.47 75.89
LensGNN-[GAT+GIN] 88.19 95.43 74.31
LensGNN-ALL 90.40 95.68 75.91
5.1. Experimental Settings
Datasets. We use three benchmark datasets whose details are given as follows. Statistics of these datasets are summarized in Table 1.
Cora is a citation network containing 2,708 research papers. These papers are categorized into seven classes. In this graph, each node represents a scientific paper, and each edge represents a citation. Bag-of-words representation is used as the feature vector for each node, which is also assigned a label indicating its research field. The task for this dataset is node classification.
PubMed is a citation network dataset containing 19,717 research papers from the biomedical field. These papers are categorized into three classes, and each paper is represented by a bag-of-words feature vector. In this graph, each node represents a scientific paper, and each edge represents a citation. The task associated with this dataset is node classification.
ogbn-arXiv is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv. In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships. These papers are categorized into 40 subject areas. We construct node attributes using the title and abstract of the papers. The task associated with this dataset is node classification.
Baselines. We compare LensGNN with 17 baselines, which can be categorized into four types. The first type involves (1) MLP. The second type involves (2) GCN (Kipf and Welling, 2016), (3) GAT (Veličković et al., 2017), (4) GIN (Xu et al., 2019), (5) GraphSAGE (Hamilton et al., 2017a), (6) Graphormer (Ying et al., 2021), (7) GT (Dwivedi and Bresson, 2012) and (8) CoarFormer (Kuang et al., 2024) seven popular GNNs for graph tasks. For fairness, we use the node representations obtained from a pre-trained LM as input to the GNN models for learning node classification. The third type involves (9) BERT (Devlin, 2018), (10) SentenceBERT (Reimers, 2019) and (11) DeBERTa (He et al., 2020) three language models. They are directly trained on graph datasets with the raw text. The fourth type involves (12) DGTL (Qin et al., 2024), (13) SNS-GPT4 (Li et al., 2024), (14) InstructGraph-INS (Wang et al., 2024), (15) InstructGLM (Ye et al., 2024), (16) GAugLLM (Fang et al., 2024) and (17) Baichuan2 (Yang et al., 2023) six LLM-based methods designed for graph tasks. We report the results from the original papers due to limitations in computational power. For Baichuan2, we directly used it for node class prediction without fine-tuning. Due to the space limitation, we move details on baselines to Appendix A.
Setup. We implement and conduct experiments for LensGNN by PyTorch. We primarily use Baichuan2-13B (Yang et al., 2023) as the backbone LLM for LensGNN. LensGNN ensemble multiple GNN models, including GCN (Kipf and Welling, 2016), GAT (Veličković et al., 2017), GIN (Xu et al., 2019), each with two layers. We perform a grid search to fine-tune hyperparameters based on the results on the validation set. The specific hyperparameter search space is shown in the Table 2. Some fixed LoRA settings included capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step. We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1. Specifically, we utilize LoRA+ (Hayou et al., 2024). For baselines that reported results on the datasets, we directly report the results from their original papers in the table. Due to computational power limitations, for those datasets where the original papers of the baselines did not report results, we left them blank in the table. We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
Table 4. Ablation study including ‘GNN Encoder’, ‘Alignment’, ‘With Text’ and ‘With Neighbor’.
[…]
These elements enhance the model’s ability to integrate representations and contextual information, resulting in higher accuracy across various datasets. GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%) - - Yes 1 87.82 93.77 74.47 GCN - Yes 1 89.29 94.23 74.35 GAT - Yes 1 90.03 94.26 73.37 GIN - Yes 1 88.92 94.82 73.26 GCN, GAT, GIN Yes No 1 19.92 40.06 34.92 GCN, GAT, GIN No Yes 1 90.77 93.96 73.73 GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03 GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
[…]
Model Bert-base T5-base(Encoder only) T5-base Falcon-7b InternLM2.5-7b-chat Baichuan2-13B
Parameters 110M 110M 220M 7B 7B 13B
Cora With GCN 81.48 78.52 76.10 85.23 89.29 89.29
With GAT 81.85 79.26 77.94 80.44 89.66 90.03
With GIN 87.50 88.24 88.24 87.45 89.29 88.92
Ensemble All 80.37 80.00 77.57 89.66 90.03 90.40
Pubmed With GCN 94.67 94.58 94.03 94.67 94.72 94.23
With GAT 94.37 94.53 94.19 94.37 95.13 94.26
With GIN 94.01 94.17 94.00 94.52 94.87 94.82
Ensemble All 94.17 93.51 93.95 94.97 95.13 95.68
5.2. Overall Performance Comparison (RQ1)
We conduct experiments to compare LensGNN with 17 state-of-the-art competitors on Cora, PubMed and ogbn-arXiv datasets. For Cora and PubMed, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set. For ogbn-arXiv, we split the dataset as suggested in (Hu et al., 2020). We use pre-trained LM to encode node representations based on their original text for all GNN models. For fairness, we use the same pre-trained LM, SentenceBERT (Reimers, 2019), as in LensGNN. Additionally, we compare our backbone LLMs, Baichuan2 (Yang et al., 2023), on the node classification task, where we use Baichuan2 for node label inference directly without fine-tuning. For LensGNN, we report the results of several variants of LensGNN , including a model that integrates GCN, GAT, and GIN simultaneously, as well as variants that ensemble each pairwise combination of these three GNNs.
Table 3 shows the classification accuracy of all the methods. From the tables, we make the following observations: (1) LensGNN outperforms the competitive methods over all the datasets. (2) Under the condition of using the same pre-trained LM to encode node representations, LensGNN outperforms individual GNN models’ classification performance across all datasets. Particularly on the PubMed dataset, LensGNN’s classification accuracy surpasses that of the best GNN methods by more than 5%. This demonstrates the effectiveness of using LLMs to ensemble multiple GNN models in node classification tasks. (3) The LM-based methods encode node representations using raw text to achieve classification. However, these approaches discard the structural properties unique to graph data, resulting in the loss of valuable information and an inability to achieve optimal performance. This indicates the necessity of structural information in graph-related tasks. (4) Some methods attempt to use LLMs to accomplish various graph tasks. For example, InstructGLM describes graph structures in natural language and feeds them into a LLM to achieve node classification. However, LensGNN still achieved better performance across all datasets because LensGNN can leverage the LLM to ensemble the strength of multiple GNNs. (5) Baichuan2 showed poor performance and could not even outperform the language models. This is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Overall, LensGNN harnesses the strengths of multiple GNNs through LLMs and avoids the tedious effort of cherry-picking GNNs across datasets, thereby achieving the best performance across all datasets.
5.3. Ablation Study (RQ2)
Figure 4. Inference Speed and Accuracy of Our Method Based on Different Backbone LLMs.
In this section, we conduct an ablation study on the Cora, PubMed, and ogbn-arXiv datasets to evaluate the impact of each model component. Specifically, we explore various configurations of the ‘GNN Encoder’ ( using different GNNs ), ‘Alignment’ ( whether multiple GNNs are aligned when training GNNs ), ‘With Text’ ( whether node text is included in prompt) and ‘With Neighbor’ ( the number of neighbour hops of the node text when inputting LLMs ). The results are presented in Table 4. From the table, we have the following observations:
Performance of different GNN encoders. We find that ensembling multiple GNNs noticeably outperforms using a single GNN. For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%. This clearly illustrates that by integrating the strengths of various types of GNNs, LensGNN can effectively enhance overall performance.
Impact of Alignment. On the PubMed dataset, the results for aligned GNN and unaligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively. This indicates that aligning multiple GNNs representation space can help LLMs better understand graph tokens. Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
The Importance of Node Text. It can be observed that node text plays a key role in the supervised fine-tuning of LLMs. Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, enabling LLM grasp structural information more deeply, thereby improving classification performance.
Overall, the ablation study confirms that each component of the model contributes significantly to its overall performance.
5.4. Performance with different LLMs (RQ3)
To further demonstrate the effectiveness of LensGNN, we use other LLMs as backbones, including three smaller LMs: BERT-base (Devlin, 2018), T5-base (Raffel et al., 2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B (Almazrouei et al., 2023) and InternLM2.5-7B-chat (Cai et al., 2024). The experiments based on these LLMs include aligning each language model separately with the representation of each type of GNN, as well as aligning them simultaneously with all three types of GNNs (GCN, GAT, GIN) and performing an ensemble. All results are presented in Table 5.
From the table, we make the following observations: (1) Smaller LMs BERT-base, T5-base, and T5-base (Encoder only) can achieve good results when ensemble a single GNN, but the performance drops significantly when ensemble multiple GNN models. This suggests that models with a smaller parameters lack the capability to integrate multiple GNNs. (2) In contrast, LLMs with large parameters that ensemble a single GNN generally perform better than smaller LMs in most cases. (3) Across different large parameter language models, the ensemble of multiple GNN models achieved the best performance on all datasets. Overall, all the observations indicate that LLMs can better achieve the ensemble of GNN models, and the multi-GNNs ensemble can harnesses the strengths of multiple GNNs to achieve the best predictions.
5.5. Model Efficiency Study (RQ4)
In this section, we analyze the computational efficiency of LensGNN during both the training and inference phases.
Model Training Efficiency. We divide the process of LensGNN into two phases: Aligning multi-GNNs and Ensembling multi-GNNs with LLM. In the first phase, the training efficiency primarily depends on the efficiency of the GNN encoders utilized, while also being related to the hyperparameter Graph token $t$. In the second phase, the training efficiency predominantly relies on the fine-tuning efficiency of the LLMs. We employ LoRA technology to reduce the number of parameters requiring fine-tuning. For instance, in the case of Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters. This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment, enabling a training batch size of 4 per device.
Model Inference Efficiency. During inference, all parameters are frozen. Since the model parameters in the first stage are much smaller than the LLM in the second stage, the time cost in the inference phase mainly comes from the second stage’s LLM inference. In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Cora and Pubmed datasets are illustrated in Figure 4. Overall, the inference speed ranges from 2 to 6 samples per second. Although a higher parameter count in the LLM results in reduced inference speed, it concurrently yields superior accuracy. In practical deployments, the choice of backbone LLM can be made by considering the balance between computational power and application requirements.
5.6. Hyperparameter Sensitivity Analysis (RQ5)
The number of Graph Tokens $t$ is an important hyperparameter, representing how many tokens each node is divided into before being input into LLM. We perform experiments on the Cora and Pubmed dataset with $t$ set to values of 1, 2, 4, 8, 16 and 32, and the results are shown in Figure 5.
The results indicate that as the number of Graph Tokens increases, the model’s performance initially rises before gradually declining. The optimal experimental results are achieved when the number of Graph Tokens is set to 8 on both datasets. When the number of tokens is too small, the distinction between nodes becomes insufficient, making it difficult for LLMs to fully learn the semantic information. On the other hand, when there are too many tokens, the representation of Graph Tokens may introduce noise or redundant information, which interferes with the model’s understanding of the graph representations. Therefore, an appropriate number of tokens helps LLMs effectively capture more complicated structural information, enhancing semantic comprehension and leading to better performance.
[…]
6. CONCLUSION
We studied the problem of how to ensemble multi-GNNs in this paper and proposed LensGNN, which ensembles multi-GNNs with LLMs. LensGNN adopts two phases of alignment: the multi-GNN alignment aims to map node representations from different GNNs into the same low-dimensional space, while the GNN-LLM alignment injects graph representation as graph tokens into LLM. After that, we perform supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology. Finally, we conducted extensive experiments to show the effectiveness of our ensembling model LensGNN. In particular, our experimental results showed that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
Appendix A Baselines
To evaluate the effectiveness of LensGNN, we compare it with the SOTA methods. Details of these baselines are summarized as follows.
(1) MLP: Multilayer Perceptron (MLP) is a type of artificial neural network that consists of multiple layers of nodes (neurons) connected by weights and is primarily used for supervised learning tasks, such as classification.
(2) Graph Neural Networks: GCN (Kipf and Welling, 2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data. GAT (Veličković et al., 2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN (Xu et al., 2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. GraphSAGE (Hamilton et al., 2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data. Graphormer (Ying et al., 2021) is built upon the standard Transformer (Vaswani, 2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks. Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data. GT (Dwivedi and Bresson, 2012) presented a simple yet effective approach to generalize transformer networks on arbitrary graphs. It includes an elegant way to fuse node positional features using Laplacian eigenvectors for graph datasets, inspired from the heavy usage of positional encodings in NLP transformer models and recent research on node positional features in GNNs. CoarFormer (Kuang et al., 2024) is a two-view architecture that captures fine-grained local information using a GNN-based module on the original graph and coarse yet long-range information using a Transformer-based module on the coarse graph.
(3) Language Models: BERT (Devlin, 2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT (Reimers, 2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa (He et al., 2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
(4) LLM Based: DGTL (Qin et al., 2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. SNS-GPT4 (Li et al., 2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. InstructGraph-INS (Wang et al., 2024) is an advanced framework that integrates instruction-based learning with graph neural networks (GNNs) to improve performance in various graph-based tasks. The model leverages both explicit instructions and implicit graph structures to enhance representation learning and decision-making processes. InstructGLM (Ye et al., 2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. GAugLLM (Fang et al., 2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. Baichuan2 (Yang et al., 2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment.
Appendix B GNN Encoders
Graph Convolutional Networks (GCNs) GCNs extend the concept of convolutional neural networks (CNNs) to graph data. They operate by aggregating and transforming feature information from a node’s local neighborhood, effectively capturing the graph’s structural information. The propagation rule in GCNs typically involves normalizing the adjacency matrix to ensure proper scaling of node features. GCNs have been widely used in tasks such as node classification, link prediction, and graph classification due to their ability to learn representations that incorporate both node features and graph topology.
Graph Attention Networks (GATs) GATs introduce an attention mechanism to the graph convolutional framework, allowing the model to learn the relative importance of neighboring nodes when aggregating their features. This is achieved by computing attention coefficients for each pair of connected nodes and using these coefficients to perform a weighted aggregation of the neighbor features. The attention mechanism in GATs makes them more flexible and capable of handling graphs with varying degrees of connectivity and node importance. GATs have shown strong performance in various graph-based tasks, providing a way to dynamically adjust the focus on relevant neighbors.
Graph Isomorphism Networks (GINs) GINs are designed to address the limitations of previous GNN models in distinguishing graph structures. They aim to be as powerful as the Weisfeiler-Lehman graph isomorphism test, which is a method for determining whether two graphs are isomorphic. GINs use a simple yet effective aggregation function that sums the features of a node’s neighbors and applies a MLP to the result. This approach ensures that GINs can capture complex graph structures and provide highly discriminative node and graph embeddings. GINs are particularly effective in graph classification tasks, where capturing subtle differences in graph topology is crucial.
Appendix C Detailed hyperparameters
The detailed hyperparameters of the first and second stages of LensGNN  are shown in Tables 6 and 7, respectively.
Table 6. Detailed hyperparameters of aligning multi-GNNs.
Model GCN GAT GIN
Dataset Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv
Hidden size 256 256 256 256 256 256 256 256 256
GNN Representation size 40960 40960 20480 40960 40960 20480 40960 40960 20480
Number of layers 2 2 2 2 2 2 2 2 2
Learning rate of GNN encoder 0.0002 0.0002 0.01 0.0002 0.0002 0.01 0.0001 0.0001 0.004
Initial learning rate of the classifier 0.0002 0.0002 0.0005 0.0002 0.0002 0.0005 0.0001 0.0001 0.0005
Dropout Rate 0.25 0.25 0.2 0.25 0.25 0.2 0.25 0.25 0.2
Weight Decay 0.25 0.25 0 0.25 0.25 0 0.25 0.25 0
Table 7. Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
Model Baichuan2-13B Falcon-7B InternLM2.5-7B-chat
Dataset Cora Pubmed ogb-arxiv Cora Pubmed Cora Pubmed
Lora dropout 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Loraplus Ir ratio 16 16 16 16 16 16 16
Prompt cutoff length 2047 2047 2047 2047 2047 2047 2047
Training batch size per device 4 4 4 4 4 4 4
Gradient accumulation steps 1 1 1 1 1 1 1
Learning rate 0.0001 0.00005 0.0001 0.0001 0.0001 0.0001 0.0001
Early stop epoch 3 2 3 3 2 3 2
Warmup Ratio 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Parameter precision FP16 FP16 FP16 FP16 FP16 FP16 FP16