### v2

# Can Large Language Models Act as Ensembler for Multi-GNNs?

Abstract
Graph Neural Networks (GNNs) have emerged as powerful models for learning from graph-structured data. However, GNNs lack the inherent semantic understanding capability of rich textual node attributes, limiting their effectiveness in applications. On the other hand, we empirically observe that for existing GNN models, no one can consistently outperforms others across diverse datasets. In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. The model first aligns multiple GNNs, mapping the representations of different GNNs into the same space. Then, through LoRA fine-tuning, it aligns the space between the GNN and the LLM, injecting graph tokens and textual information into LLMs. This allows LensGNN to ensemble multiple GNNs and take advantage of the strengths of LLM, leading to a deeper understanding of both textual semantic information and graph structural information. The experimental results show that LensGNN  outperforms existing models. This research advances text-attributed graph ensemble learning by providing a robust and superior solution for integrating semantic and structural information. We provide our code and data here: https://anonymous.4open.science/r/EnsemGNN-E267/.
Can Large Language Models Act as Ensembler for Multi-GNNs?
Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
1 INTRODUCTION
Graphs are structured data that captures the inter-relations between entities in the real world. To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced Kipf and Welling (2016) and message passing is the main mechanism for neighborhood aggregation Hamilton et al. (2017a); Veličković et al. (2017). GNNs have demonstrated significant success across various applications such as social network analysis Hamilton et al. (2017a), recommendation systems Wang et al. (2019), and molecular property prediction Gilmer et al. (2017).
Figure 1: Comparison between GNN models on node classification across different datasets.
Despite the success, two major challenges remain unresolved. First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes Zhao et al. (2019). This can lead to the loss of valuable semantic information during the learning process, limiting the effectiveness of GNNs in applications where node features contain meaningful texts, such as citation networks and social media platforms. Second, while there have been a series of GNNs proposed, no one has been shown to consistently lead others across datasets of various domains. For example, we compare the node classification performance of four representative GNNs, namely, APPNP Gasteiger et al. (2018), GAT Veličković et al. (2017), GCN Kipf and Welling (2016) and GIN Xu et al. (2019) on three benchmark datasets Yan et al. (2023): Cora, Citeseer and PubMed. For fairness, we set the equal number of hidden representation layers and also the same dimensionality size. The results given in Figure 1 demonstrate that the winner varies across the datasets. The challenge of selecting the optimal GNN for a given dataset remains unresolved, as different GNN architectures exhibit varying strengths. This also restricts the wide applicability of GNNs.
To remedy the incapability of GNNs in semantic understanding, existing works Qin et al. (2024); Li et al. (2024); Wang et al. (2024); Ye et al. (2024); Fang et al. (2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation. The inherent advantage of GNNs in utilizing graph structure, fused with the strength of LLMs in semantic understanding, mutually enhances both models and leads to superior performance in downstream tasks. However, how to select the optimal GNN in different scenarios remains a gap. Since the effectiveness of GNNs varies across datasets, there naturally arises a question: Can we develop an ensembled model for multi-GNNs? The ensembled model is expected to integrate the strengths of multiple GNNs and can consistently perform well across datasets. To further leverage the power of LLMs in text understanding, we thus upgrade the question: Can LLMs act as ensembler for multi-GNNs?
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
  • We propose a novel method LensGNN to ensemble multi-GNNs with LLMs.
  • We introduce an effective graph prompt tuning paradigm with both the multi-GNN alignment and the GNN-LLM alignment.
  • We conduct extensive experiments to show the superiority of LensGNN.
2 RELATED WORK
GNNs Kipf and Welling (2016); Xu et al. (2019); Veličković et al. (2017); Hamilton et al. (2017b) have been widely used in learning from graph-structured data. Despite their advantages, GNNs struggle to process textual attributes effectively, as traditional GNNs lack the semantic understanding capability that is necessary to handle text-attributed graphs (TAGs). In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) Zoph and Le (2017). However, it is very computationally expensive, which necessitates new exploration.
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al. (2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, like GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024) GraphLLM Chai et al. (2023) and ENGINE Zhu et al. (2024). Another notable category is to employ LLMs as encoders, as seen in the works of OFA Liu et al. (2024), TextGNN Zhu et al. (2021) and AdsGNN Li et al. (2021), which focus on optimization strategies for encoding graph data. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods, such as GOFA Kong et al. (2024), LTRN Zhang et al. (2021) and GLEM Zhao et al. (2022). These diverse approaches highlight the evolving landscape of LLM and GNN integration. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
3 PRELIMINARIES
Text-attributed graph. A text-attributed graph is a graph where nodes and edges have textual attributes. In this work, we only consider the case where only nodes have textual attributes. Formally, a text-attributed graph can be defined as $G=(V,E,X^{(v)})$, where $V=\{v_{1},v_{2},\ldots,v_{|V|}\}$ represents the set of nodes, and $E=\{e_{1},e_{2},\ldots,e_{|E|}\}$ represents the set of edges, with $|V|=N$ indicating the total number of nodes in the graph. $X^{(v)}=\{x^{(v)}_{1},x^{(v)}_{2},\ldots,x^{(v)}_{|V|}\}$ denote the attributes on nodes, respectively, which are strings. It can be represented as $G=(V,E,\{x_{n}\}_{n\in V})$. Additionally, we define $A$ as the adjacency matrix of graph $G$, where the matrix size is $N\times N$. In an unweighted graph, $A(i,j)=1$ indicates that there is an edge between nodes $v_{i}$ and $v_{j}$, while $A(i,j)=0$ indicates that there is no edge between them.
Graph Neural Networks. Traditional GNNs are a class of deep learning models designed for handling graph-structured data. The basic architecture of a GNN includes a node representation layer and a message-passing layer. In the node representation layer, each node $v$ is assigned a feature vector $x_{v}$. In the message-passing layer, the representation vector $h_{v}^{(t)}$ of node $v$ is updated after the $t$-th iteration using the following formula:
$h_{v}^{(t)}=\text{UPDATE}\left(\text{AGG}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$ (1)
where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$. The AGG function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state. Through this iterative process, GNNs are able to learn increasingly rich node representations, capturing both local and global structural information in the graph. The specific implementations of the AGG and UPDATE functions can vary depending on the particular GNN model.
4 METHODOLOGY
This section introduces the main steps in LensGNN, which includes aligning multi-GNNs and ensembling multi-GNNs with LLM. The overall model procedure is given in Fig. 2.
4.1 Aligning multi-GNNs
Given multiple GNNs, they could generate node representations in different low-dimensional spaces. Therefore, before ensembling GNNs, we need to first align them.
Figure 2: The overall process of LensGNN.
Node feature initialization. To utilize the rich semantic information of textual features, we use pre-trained language models to initialize them. Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT Reimers (2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
$\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$ (2)
This process allows texts of any dataset and length to be converted into the same feature space.
GNN representations. Next, the model takes the text representations obtained from Eq. 2 for each node and feeds them into multiple GNNs. For the $k$-th GNN, we denote the node embedding matrix $H^{[k]}$ generated by the final layer as:
$H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$ (3)
where $\tilde{X}$ is the initial node embedding matrix from Eq. 2 and $\Theta_{G_{k}}$ is the learnable parameters of the $k$-th GNN. For each node $v_{i}$, let $h_{i}^{[k]}$ be its embedding vector, which is the $i$-th row in $H^{[k]}$.
Reshape node representations. The output of each GNN is then connected to a linear layer, which transforms node representations into the dimensionality of hidden embeddings in LLMs. This step paves the way for the subsequent alignment of GNN and LLM. Details will be given in the next section. For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$ (4)
Note that the shape of $\tilde{h}^{[k]}_{i}$ is a one-dimensional vector of length $t\times e$, where $t$ is a hyperparameter indicating the number of graph tokens each node representation is mapped to. The dimensionality $e$ comes from the hidden embedding layer of the LLM used, where each token is mapped to an embedding of shape $(1,e)$.
Multi-GNN alignment. To align node representations from multiple GNNs, we next feed them into a shared linear layer, which serves as the classifier. During the training time, we alternatively input representations from different GNNs into the classifier and train them sequentially. This training approach allows node representations from multi-GNNs to be aligned, which facilitates the integration of multi-GNNs. After training, the parameters of GNNs and the linear layer in Eq. 4 will be frozen, which are then used in aligning GNNs and LLMs. The overall procedure for GNN alignment is summarized in the top of Fig. 2.
In summary, during the multi-GNN alignment step, we train each GNN to ensure that their outputs reside in the same vector space. By incorporating the node feature initialization step that utilizes language pre-trained models, our model gains the ability to extract semantic and structural information from the nodes in the text graph. Building on the full potential of GNNs, we map the dimensions of the GNN representations to the dimensions required by the LLM embeddings, thereby maximizing the integration of extensive graph structural information into the LLM in the subsequent steps.
4.2 Ensembling multi-GNNs with LLM
After GNNs are trained, we can get node representations capturing rich graph structural information. Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure Guo et al. (2023). Therefore, in the second stage, we empower LLMs to comprehend graph representations. Further, with specially designed prompts, we enable these models to implicitly perform ensembling.
Align GNNs and LLMs. The key step in enabling LLMs to understand GNN tokens lies in aligning GNN representations, which encapsulate rich structural information, with the semantic space required by LLMs. This process necessitates fine-tuning LLMs. However, due to the extensive number of parameters in these models, fine-tuning requires significant computational resources, which is challenging for practical applications. As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector Tang et al. (2024), mapping GNN representations into the semantic space. While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
[…]
Prompt design and Multi-GNN ensembling. After aligning graph structure and semantics, LLM can understand graph representations. We next design prompts and take LLM as the ensembler to integrate multi-GNNs. Our goal is to leverage the strengths of both multi-GNNs and LLM to output more accurate predictions. The prompt design follows three key principles:
  1. 1.
    Simultaneous input of textual strings from target node and its adjacent neighbors, graph tokens of the node, and task specification. Existing studies Kipf and Welling (2016); Veličković et al. (2017); Verma et al. (2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction. Further, clear task instructions are necessary for accurate prediction.
  2. 2.
    Differentiation between text tokens and graph tokens. It directs LLM to correctly distinguish between textual tokens and graph tokens within different segments of the prompt, thus preventing confusion.
  3. 3.
    Guidance for learning from multi-GNNs: It leads LLM to implicitly learn how to effectively combine strengths of different GNNs.
Based on the above principles, we provide a prompt template used in our experiments. Details on the prompt template are given in Appendix D. Further, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
5 EXPERIMENT
Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods. LensGNN utilizes Baichuan2-13B-chat as the backbone LLM. We highlight the best score on each dataset in bold and underline the runner-up’s.
Model type Model Cora PubMed ogbn-arXiv Citeseer Wiki-CS
MLP MLP 66.42 82.40 61.47 71.13 68.41
GNN GCN 85.97 83.78 70.81 77.74 77.15
GAT 86.71 83.59 70.05 78.21 78.94
GIN 85.60 82.03 67.03 75.07 68.97
GraphSAGE 84.87 87.01 70.35 73.61 79.56
Graphormer 80.41 88.75 72.81 71.28 72.07
Language Model BERT 80.15 93.91 72.78 73.17 78.33
SentenceBERT 78.82 92.49 71.42 72.79 77.92
DeBERTa 77.79 93.45 72.90 73.13 75.11
LLM Ensembler LensGNN-[GCN+GAT] 88.56 94.11 75.78 78.05 18.41
LensGNN-[GCN+GIN] 91.88 94.47 75.89 76.64 83.47
LensGNN-[GAT+GIN] 88.19 95.43 74.31 78.99 82.98
LensGNN-ALL 90.40 95.68 75.91 79.31 81.78
5.1 Experimental settings
Datasets. We use eight benchmark datasets, comprising five node classification datasets: Cora, PubMed, ogbn-arXiv, Citeseer and Wiki-CS, and three molecular graph classification datasets Wu et al. (2018): BACE, BBBP and ClinTox. Details on these datasets are provided in the Appendix A.
Baselines. We compare our method with 19 baselines: MLP, GNN models, LM-based models and LLM-based models. Specifically, GNN models include GCN Kipf and Welling (2016), GAT Veličković et al. (2017), GIN Xu et al. (2019), GraphSAGE Hamilton et al. (2017a) and Graphormer Ying et al. (2021). LM-based models consist of BERT Devlin (2018), SentenceBERT Reimers (2019) and DeBERTa He et al. (2020). LLM-based models include DGTL Qin et al. (2024), SNS-GPT4 Li et al. (2024), GAugLLM Fang et al. (2024) and Baichuan2-13B Yang et al. (2023), which all employ 13B LLMs. We also compare LensGNN with other 7B-LLM-based models: GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024), OFA Liu et al. (2024) and GOFA Kong et al. (2024). Due to the space limitation, details on these baselines are given in Appendix B.
Setup. We use Baichuan2 Yang et al. (2023) and InternLM2.5 Cai et al. (2024) as the backbone LLM for LensGNN. In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers. For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input. For more details on experiment setup, refer to Appendix C.
5.2 Node classification results
Since the large number of baselines report their results on different datasets and settings, we show the classification results in three tables, respectively. Table 1 compares LensGNN with MLP, GNN models and small language models. Tables 2 summarizes the results against LLM-based models. Table 3 evaluates the performance of LensGNN in the 20-shot setting. From these tables, we observe:
(1) LensGNN consistently outperforms all the GNN models and LM-based methods across all the datasets. For GNN models, while they initialize node features with SentenceBERT to capture text semantics, LLMs have much stronger capability in text understanding and generation, which explains the advantage of LensGNN. For LM-based models, in addition to the weak expressiveness of small language models, they ignore the rich structural information of graphs, degrading their performance.
(2) LensGNN achieves better performance than LLM-based models, even in the 20-shot setting. In particular, we see that Baichuan2-13B performs very poorly. We speculate this is because LLMs that have not been fine-tuned are not suitable for node classification tasks on graph data. Different from other LLM-based models, LensGNN integrates the strengths of multiple GNNs and employs two-phases of alignment, which explains its superior performance.
Table 2: Comparison on classification accuracy (%) with LLM-based models. * denotes the best variant of our model and - means that the results are missing from their original papers.
Model Cora PubMed ogbn-arXiv
GraphGPT-MIX-7B - 74.16 64.76
LLaGA-HO-7B(GENERAL) - 94.45 75.01
InstructGLM-Llama-7B 87.08 93.84 75.70
DGTL 81.10 87.10 -
SNS-GPT4 82.50 93.80 74.40
GAugLLM - 83.68 74.15
Baichuan2-13B 13.65 36.04 4.79
LensGNN*-InternLM2.5-7B 90.03 95.13 74.24
LensGNN*-Baichuan2-7B 89.85 95.08 75.88
LensGNN*-Baichuan2-13B 91.88 95.68 75.91
Table 3: Classification accuracy in 20-shot setting.
Model Cora PubMed
GCN 73.5 68.0
GAT 72.8 68.1
GIN 75.7 69.3
OFA 75.34 77.89
GOFA 77.08 87.33
LensGNN*-Baichuan2-7B 78.09 89.19
LensGNN*-InternLM2.5-7B 79.93 87.75
5.3 Graph classification results
Table 4: Graph classification results.
Model BACE BBBP ClinTox
ACC AUC ACC AUC ACC AUC
GCN 74.01 84.92 84.14 79.00 90.93 87.23
GAT 74.67 80.85 76.34 66.12 91.61 86.37
GIN 68.09 85.66 83.65 79.02 86.91 85.80
LensGNN-[GCN+GAT] 73.02 72.67 86.09 81.29 98.99 94.00
LensGNN-[GCN+GIN] 75.32 74.59 88.04 83.27 99.32 96.00
LensGNN-[GAT+GIN] 76.31 75.99 86.09 81.64 98.99 95.81
LensGNN-ALL 80.59 80.33 88.53 85.37 98.99 97.63
We further evaluate the effectiveness of LensGNN on the graph classification task. We use classification accuracy (ACC) and AUC scores as the evaluation metrics. The results are given in Table  4. From the table, LensGNN outperforms GNN baselines in most cases, which shows that combining GNNs and LLMs is inspiring. Further, LensGNN-ALL generally performs better than its variants. This verifies that leveraging strengths of multiple GNNs is useful. Surprisingly, we find that on BACE, LensGNN and its variants achieve smaller AUC scores than GNN models. We step into the datasets and observe that BACE is a label-balanced dataset while BBBP and ClinTox are highly imbalanced. Note that, for balanced dataset, ACC is more convincing than AUC score, while AUC score is a better indicator than ACC on imbalanced dataset. Therefore, the ACC advantage of LensGNN on BACE and the AUC leads on other two datasets have evidently demonstrated the superiority of our method.
5.4 Ablation study
We next systematically conduct an ablation study on Cora, PubMed, and ogbn-arXiv datasets to evaluate the importance of major model components. Specifically, we explore various configurations of “GNN Encoder” (using different GNNs ), “Alignment” (whether multiple GNNs are aligned when training GNNs ), “With Text” (whether node text is included in prompt) and “With Neighbor” (whether neighborhood information is included or not). The results are presented in Table 5.
Performance of different GNN encoders. Ensembling multiple GNNs noticeably outperforms using a single GNN. For example, on the ogbn-arxiv dataset, the best result from a single GNN is 74.35%, while LensGNN achieves 75.91%. This illustrates that by integrating the strengths of various GNNs, LensGNN can effectively enhance the overall performance.
Impact of Alignment. On the PubMed dataset, the results for unaligned GNN and aligned GNN are 93.96% and 95.68%, respectively, while on the ogbn-arXiv dataset, the results are 73.73% and 75.91%, respectively. This shows that multi-GNN alignment can help LLMs better understand graph tokens. Although the performance of aligned GNN is slightly lower than unaligned GNN on the Cora dataset, this may be due to the insufficient number of training samples.
Importance of Node Text. It can be observed that node text plays a key role in the supervised fine-tuning of LLMs. Through these node texts, the model can achieve a deeper understanding of semantics, resulting in outstanding accuracy.
Role of Neighbors. The textual information of neighbors provides support for LLM’s understanding of the semantic information of nodes, thereby improving classification performance.
Table 5: Ablation study results.
GNN Encoder Alignment With Text With Neighbor Cora (Acc/%) PubMed (Acc/%) ogbn-arXiv (Acc/%)
- - Yes 1 87.82 93.77 74.47
GCN - Yes 1 89.29 94.23 74.35
GAT - Yes 1 90.03 94.26 73.37
GIN - Yes 1 88.92 94.82 73.26
GCN, GAT, GIN Yes No 1 19.92 40.06 34.92
GCN, GAT, GIN No Yes 1 90.77 93.96 73.73
GCN, GAT, GIN Yes Yes 0 84.87 94.21 72.03
GCN, GAT, GIN Yes Yes 1 90.40 95.68 75.91
5.5 Performance with different backbones
Table 6: The classification accuracy (%) with different LLMs on Cora and PubMed.
Model Bert-base T5-base (Encoder only) T5-base Falcon-7B InternLM2.5-7B-chat Baichuan2-13B
Parameters 110M 110M 220M 7B 7B 13B
Cora With GCN 81.48 78.52 76.10 85.23 89.29 89.29
With GAT 81.85 79.26 77.94 80.44 89.66 90.03
With GIN 87.50 88.24 88.24 87.45 89.29 88.92
Ensemble All 80.37 80.00 77.57 89.66 90.03 90.40
Pubmed With GCN 94.67 94.58 94.03 94.67 94.72 94.23
With GAT 94.37 94.53 94.19 94.37 95.13 94.26
With GIN 94.01 94.17 94.00 94.52 94.87 94.82
Ensemble All 94.17 93.51 93.95 94.97 95.13 95.68
To study the impact of backbone models on LensGNN, we use different backbones, including three small LMs: BERT-base Devlin (2018), T5-base Raffel et al. (2023), the encoder-only variant of T5-base, and two other LLMs: Falcon-7B Almazrouei et al. (2023) and InternLM2.5-7B-chat Cai et al. (2024). All the results are presented in Table 6. From the table, we see that
(1) Small LMs BERT-base, T5-base, and T5-base (Encoder only) perform well when applied to a single GNN, but the performance drops significantly when ensembling multiple GNN models. This suggests that models with less parameters lack the capability to integrate multiple GNNs.
(2) For LLMs, they generally perform better than LM backbones. For example, for GAT on Cora, the best result for LM backbones is 81.85%, while that for LLM backbones is 90.03%. This is because LLMs can provide richer text semantics. Further, compared with single GNN, the ensemble of multiple GNNs with LLMs leads to better performance, which verifies that unifying the strengths of multiple GNNs is beneficial for node classification.
5.6 Model efficiency study
Training Efficiency. The training of LensGNN includes two main phases: aligning multi-GNNs and ensembling multi-GNNs with LLM. In the first phase, the time complexity depends on the GNN encoders. Since the used GNNs have a linear time complexity w.r.t. the number of nodes in the graph, the training process is efficient. In the second phase, fine-tuning LLMs is the main computational cost. We employ LoRA to reduce the number of parameters to be fine-tuned. For example, for Baichuan2-13B, the trainable parameters constitute only 0.0470% of the total parameters. This allows us to fine-tune the 13B parameter LLM in a single 80G Nvidia A100 environment. Compared to GraphGPT Tang et al. (2024), the increase of our training computational cost primarily stems from employing more than one GNN. However, in contrast to LLM, the number of parameters in GNN is considerably smaller, which increases marginal training cost.
Inference Efficiency. During inference, all the parameters are frozen. The major cost of model inference comes from LLMs. In our experiments, with a maximum input length of 2047 tokens, the inference speed and accuracy of our method based on Baichuan2-13B, Falcon-7B, and InternLM2.5-7B-chat on the Pubmed dataset is illustrated in Figure 3. Overall, the inference speed ranges from $2$ to $4$ samples per second. Although a larger LLM results in reduced inference speed, it concurrently yields superior accuracy.
Figure 3: Inference speed and accuracy comparison.
5.7 Hyperparameter sensitivity analysis
We end this section with model sensitivity analysis on the number of graph tokens $t$, which represents how many tokens each node is mapped into before fed into LLM. We conduct experiments on Cora and Pubmed with varying $t$ values, and the results are shown in Figure 4. From the figure, we see that, as the number of graph tokens increases, the model’s performance first rises and then decreases. For both datasets, the best results are achieved when $t=8$. When $t$ is small, graph token representations cannot well capture the semantic information. On the other hand, when $t$ is large, the representation of graph tokens could be noise-corrupted, which adversely affects the model’s understanding on graph structure. Therefore, an appropriate number of tokens helps LLMs effectively capture structural information, enhancing semantic comprehension and leading to better performance.
Figure 4: Hyperparameter sensitivity analysis.
6 CONCLUSION
We studied the problem of how to ensemble multi-GNNs in this paper and proposed LensGNN, which ensembles multi-GNNs with LLMs. LensGNN adopts two phases of alignment: the multi-GNN alignment aims to map node representations from different GNNs into the same low-dimensional space, while the GNN-LLM alignment injects graph representation as graph tokens into LLM. After that, we perform supervised fine-tuning with LoRA to enhance the LLM’s capability in understanding graph topology. Finally, we conducted extensive experiments to show the effectiveness of our ensembling model LensGNN. In particular, our experimental results showed that LLMs can serve as effective multi-GNN ensembler, while small LMs cannot.
7 LIMITATIONS
Our method could suffer from several limitations:
First, due to limited computational resources, we do not conduct experiments on LLMs with parameters exceeding 13B. This implies that the upper limit of our method’s capabilities has not been fully investigated. For larger-scale LLMs that are utilized in production environments, they could further improve our model’s performance.
Second, our method relies on manually crafted prompts to ensemble multi-GNNs. Future studies on prompt design could be a promising direction.
Third, we only evaluated the performance based on node classification and graph classification tasks. Evaluation on other tasks, such as graph dataset generation and graph task interpretability analysis, are valuable research questions.
8 ETHICAL STATEMENT
Our work falls under basic research and is not tied to specific applications; therefore, whether our method will be abused and cause negative social impacts depends on the specific applications in which others use our method. In addition, our work does not involve any stakeholders benefiting or being disadvantaged, nor does it involve vulnerable groups. The datasets we used are all commonly used public datasets, posing no privacy risks, and aligned with their intention for scientific research. For the datasets, pre-trained models, and training frameworks utilized, we adhered to their respective open-source licenses: community license for Baichuan-2 Yang et al. (2023), Apache License 2.0 for InternLM2.5 Cai et al. (2024) , Falcon-7B Almazrouei et al. (2023) and LLaMA-Factory Zheng et al. (2024)
Appendix A Datasets
We use eight benchmark datasets, comprising five node classification datasets: Cora, PubMed, ogbn-arXiv, Citeseer and Wiki-CS, and three molecular graph classification datasets: BACE, BBBP and ClinTox. Statistics of these datasets are summarized in Table 7.
Cora McCallum et al. (2000) is a standard citation network dataset consisting of 2,708 research papers in the field of machine learning. These papers are categorized into seven classes. The title and abstract of each paper are utilized as the textual attributes of the nodes. The dataset includes 5,429 citation links, constructing a graph structure among the papers. The Cora dataset is commonly used for node classification and evaluating the performance of graph neural network models.
Citeseer Giles et al. (1998) is a citation network dataset comprising 3,312 research papers primarily from the fields of computer science and information technology. These papers are categorized into six classes based on their research areas. The title and abstract of each paper are used as the textual attributes of the nodes, providing a semantic representation of the content. The dataset includes 4,732 citation links, forming a graph structure that represents the citation relationships between the papers. The Citeseer dataset is widely used for node classification tasks, and it serves as a benchmark for evaluating the performance of graph neural network models.
PubMed Sen et al. (2008) is another citation network dataset containing 19,717 research papers from the biomedical field. These papers are categorized into three classes. The title and abstract of each paper are utilized as the textual attributes of the nodes. The dataset includes 44,338 citation links, constructing a graph structure among the papers. The PubMed dataset is also used for node classification tasks, particularly in testing large-scale graph neural network models.
ogbn-arXiv Hu et al. (2020) is part of the Open Graph Benchmark (OGB) and contains 169,343 academic papers scraped from arXiv. These papers are time-ordered by their submission dates and categorized into 40 subject areas. In this dataset, the nodes represent arXiv papers, and the edges represent citation relationships. The dataset includes 1,166,243 citation links, constructing a citation network among the papers. The ogbn-arXiv dataset is commonly used for node classification and studying time-sensitive graph neural network models.
Wiki-CS Mernyei and Cangea (2022) is derived from Wikipedia and consists of 11,701 web pages related to computer science topics. These pages are categorized into 10 classes, each representing a different area of computer science such as artificial intelligence, computer architecture, and software engineering. The text content of each page is used as the textual attributes of the nodes, capturing the thematic essence of the pages. The dataset includes 216,123 hyperlinks, which construct a graph structure connecting the web pages. The Wiki-CS dataset is often utilized for node classification tasks and for evaluating graph-based machine learning models.
BACE Wu et al. (2018) is a collection of inhibitors of human beta-secretase 1 (BACE-1) and provides both quantitative IC50 values and qualitative binary labels indicating the binding results. The BACE dataset comprises 1,513 molecular graphs for the molecular property prediction.
BBBP Wu et al. (2018) is used for predicting blood-brain barrier permeability (BBBP), which is crucial for determining whether a molecule can cross the blood-brain barrier. The BBBP dataset contains 2,039 molecular graphs for the binary graph classification task.
ClinTox Wu et al. (2018) is a collection of drugs approved by the FDA and those that have failed clinical trials due to toxicity reasons. The dataset encompasses two classification tasks for 1,491 drug molecules: (1) clinical trial toxicity and (2) FDA approval status.
For Cora, PubMed, Citeseer, BACE, BBBP and ClinTox, we randomly split nodes into 60%, 20%, and 20% for training, validation and testing, and measure the performance of all models on the test set.
For ogbn-arXiv we split the dataset as suggested in Hu et al. (2020).
For Wiki-CS, we split the dataset as suggested in Mernyei and Cangea (2022).
Table 7: Datasets statistics.
Dataset #Graphs Avg.# Nodes Avg.# Edges # Classes
Cora 1 2,708 5,429 7
PubMed 1 19,717 44,338 3
ogbn-arXiv 1 169,343 1,166,243 40
Citeseer 1 3,312 4,732 6
Wiki-CS 1 11,701 216,123 10
BACE 1,513 34.1 73.7 1
BBBP 2,039 23.9 51.6 1
ClinTox 1,491 26.1 55.5 2
Appendix B Baselines
To evaluate the effectiveness of LensGNN, we compare it with the SOTA methods. Details of these baselines are summarized as follows.
(1) MLP: Multilayer Perceptron (MLP) is a type of artificial neural network that consists of multiple layers of nodes (neurons) connected by weights and is primarily used for supervised learning tasks, such as classification.
(2) Graph Neural Networks: GCN Kipf and Welling (2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data. GAT Veličković et al. (2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN Xu et al. (2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. GraphSAGE Hamilton et al. (2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data. Graphormer Ying et al. (2021) is built upon the standard Transformer Vaswani (2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks. Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data.
(3) LM Based Models: BERT Devlin (2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT Reimers (2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa He et al. (2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
Figure 5: A prompt template for node classification task.
Figure 6: A prompt template for graph classification task.
(4) LLM Based Models: GraphGPT Tang et al. (2024) is a novel framework that integrates Large Language Models (LLMs) with graph structural knowledge through graph instruction tuning. This innovative approach enables LLMs to comprehend and interpret the structural components of graphs, thereby demonstrating superior generalization in both supervised and zero-shot graph learning tasks. LLaGA Chen et al. (2024) integrates the capabilities of LLMs with graph-structured data, enabling LLMs to handle complex graph tasks effectively. LLaGA achieves this by reorganizing graph nodes into structure-aware sequences and mapping them into the token embedding space through a versatile projector, demonstrating superior versatility, generalizability, and interpretability across various datasets and tasks. InstructGLM Ye et al. (2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. DGTL Qin et al. (2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. SNS-GPT4 Li et al. (2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. GAugLLM Fang et al. (2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. Baichuan2 Yang et al. (2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment. OFA Liu et al. (2024) is a pioneering framework that unifies various graph tasks into a single model, enabling it to address classification tasks across different domains and tasks. OFA achieves this by converting graph data into text-attributed graphs (TAGs), using language models to encode diverse text attributes into a common embedding space, and introducing a novel graph prompting paradigm for in-context learning without fine-tuning. GOFA Kong et al. (2024) is a novel graph foundation model that integrates the strengths of large language models (LLMs) and graph neural networks (GNNs) to enable joint graph and language modeling. It achieves this by interleaving GNN layers into a frozen pre-trained LLM, allowing for self-supervised pretraining on graph data and fluidity in handling various graph-related tasks.
Table 8: Hyperparameter Search Spaces.
Hyperparameter Search space
learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
dropout $[0,0.25]$
loraplus lr ratio $\{16.0,24.0,32.0\}$
Graph token $t$$\{1,2,4,8,16\}$
Table 9: Detailed hyperparameters of aligning multi-GNNs.
Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
GCN Cora 256 40960 2 0.0002 0.0002 0.25 0.001
Pubmed 256 40960 2 0.0002 0.0002 0.25 0.001
ogbn-arxiv 256 20480 2 0.01 0.0005 0.2 0
Citeseer 256 40960 2 0.0001 0.0005 0.25 0.001
Wiki-CS 256 40960 2 0.0002 0.0005 0.25 0.001
BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
GAT Cora 256 40960 2 0.0002 0.0002 0.25 0.001
Pubmed 256 40960 2 0.0002 0.0002 0.25 0.001
ogbn-arxiv 256 20480 2 0.01 0.0005 0.2 0
Citeseer 256 40960 2 0.0001 0.0005 0.25 0.001
Wiki-CS 256 40960 2 0.0002 0.0005 0.25 0.001
BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
GIN Cora 256 40960 2 0.0001 0.0001 0.25 0.001
Pubmed 256 40960 2 0.0001 0.0001 0.25 0.001
ogbn-arxiv 256 20480 2 0.004 0.0005 0.2 0
Citeseer 256 40960 2 0.00005 0.0005 0.25 0.001
Wiki-CS 256 40960 2 0.0001 0.0005 0.25 0.001
BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
Table 10: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
Baichuan2-13B Cora 0.1 16 4 0.0001 3 0.1
Pubmed 0.1 16 4 0.00005 2 0.1
ogbn-arxiv 0.1 16 4 0.0001 3 0.1
Citeseer 0.15 16 4 0.0001 5 0.1
Wiki-CS 0.1 16 4 0.0001 5 0.1
BACE 0.1 16 4 0.0001 4 0.1
BBBP 0.1 16 4 0.0001 3 0.1
ClinTox 0.1 16 4 0.0001 3 0.1
Baichuan2-7B Cora 0.15 16 4 0.0001 3 0.1
Pubmed 0.15 16 4 0.0001 2 0.1
ogbn-arxiv 0.15 16 4 0.0001 3 0.1
InternLM2.5-7B Cora 0.1 16 4 0.0001 3 0.1
Pubmed 0.1 16 4 0.0001 2 0.1
ogbn-arxiv 0.1 16 4 0.0001 3 0.1
Falcon-7B Cora 0.1 16 4 0.0001 3 0.1
Pubmed 0.1 16 4 0.0001 2 0.1
Appendix C Experiment setup
We implement LensGNN by PyTorch. We primarily use Baichuan2 Yang et al. (2023) and InternLM2.5 Cai et al. (2024) as the backbone LLM for LensGNN. In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers. We perform grid search to fine-tune hyperparameters based on the validation set. Details on the search space is given in Table 8. We utilize LoRA+ Hayou et al. (2024) for fine-tuning. Some fixed LoRA settings include capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step. We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1. For the training of LLMs, we utilize LLaMA-Factory Zheng et al. (2024) as a framework. For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input. For baselines that report results on the adopted datasets, we directly report the results from their original papers. For those whose results are missing, we leave them blank. For Baichuan2, we directly use it without fine-tuning. We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
Appendix D Prompt design
The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific text of target node or graph and their multiple GNN representations. An example of a prompt template can be seen in Figure 5. For graph classification, we follow MolXPT Liu et al. (2023) to design prompts for molecular graphs as shown in Figure 6.
Appendix E Detailed hyperparameter settings
The details on the setting of hyperparameters in the first and second stages of LensGNN  are shown in Tables 9 and 10, respectively. The hyperparameters for the 20-shot experiment are shown in Tables 11 and 12.
Table 11: Detailed hyperparameters of aligning multi-GNNs (20-shot).
Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
GCN Cora 256 32768 2 0.0002 0.0005 0.25 0.001
Pubmed 256 32768 2 0.0002 0.0005 0.25 0.001
GAT Cora 256 32768 2 0.0002 0.0005 0.25 0.001
Pubmed 256 32768 2 0.0002 0.0005 0.25 0.001
GIN Cora 256 32768 2 0.0001 0.0005 0.25 0.001
Pubmed 256 32768 2 0.0001 0.0005 0.25 0.001
Table 12: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune) (20-shot).
Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
Baichuan2-7B Cora 0.15 16 4 0.0001 8 0.1
Pubmed 0.15 16 4 0.0001 9 0.1
InternLM2.5-7B Cora 0.1 16 4 0.0001 8 0.1
Pubmed 0.05 16 4 0.0001 9 0.1